How many gallons of a 12% salt solution must be mixed with 35 gallons of a 28% solution to obtain a 20% solution?
Answer by richwmiller(17219)

.28*35+.12x=.20*(x+35)
0.12x=(x+81.6667) = 0.20(x+35.)
0.12x+9.8 = 0.20x+7.
2.8=0.08x
x = 35